# Behavior-Dependent Linear Recurrent Units for Efficient Sequential Recommendation

### v1

@@ Abstract.
Recommender Systems, Sequential Recommendation, Recurrent Neural Networks
copyright: acmcopyrightjournalyear: 2018doi: 10.1145/1122445.1122456conference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Sequential recommender systems are crucial for personalized online services, capturing evolving user preferences through past interactions and predicting future actions by effectively modeling sequential dependencies within historical user behaviors (chen2018sequential; xie2022contrastive). The core of sequential recommendation is the user behavior modeling techniques, which have evolved significantly over time, advancing from traditional methods such as Markov chains (he2016fusing; rendle2010factorizing) to sophisticated deep learning methods (hidasi2015session; yan2019cosrec). As illustrated in Figure 1, there is a triangle of golden principles for designing the user behavior modeling operators for sequential recommendation, i.e., training efficiency, low-cost inference, and strong performance.
@@
We choose the linear recurrent unit (LRU) (martin2017parallelizing; orvieto2023resurrecting) as the basic operator, and start by investigating its potential drawbacks in the case of sequential recommendation scenarios when considering the triangle of three golden principles. Firstly, the LRU operator is independent to the input user behaviors, i.e., every single user behavior is transformed through the same network, thereby contributing equally to the user preference representation. However, the diverse behaviors in the user sequence should make different contributions due to the multifaceted nature of user interests. Hence, it is essential to build a behavior-independent operator to improve the performance of user behavior modeling. Secondly, previous work, i.e., LRURec (yue2024linear), simply follows the classical design of LRU by leveraging the eigenvalue decomposition in the complex domain. Nevertheless, as demonstrated in (gu2023mamba), complex elements are beneficial for continuous modalities like video and audio, but not for discrete sequence modeling like user behaviors. Moreover, complex-domain LRUs introduce a larger number of parameters compared to real-valued counterparts, potentially leading to overfitting and degenerated performance. Lastly, while the above two points both aim at the “strong performance” principle, they can largely hurt the training and inference efficiency due to the complicated design of behavior modeling operators (i.e., behavior-dependent and non-complex). Therefore, it is crucial to move beyond the simple model design, and consider hardware-aware parallelism acceleration algorithm for the rest two principles of “training efficiency” and “low-cost inference” simultaneously.
@@ 3.3. Hardware-Aware Parallel Scan
Since modern deep learning accelerators like GPUs are only optimized for operators like matrix multiplications and convolutions, we develop an efficient custom CUDA kernel for our BD-LRUs, implemented in OpenAI’s Triton (tillet2019triton), which is a language and compiler for GPU programming and parallel computing. Our hardware-aware algorithm is based on scans (blelloch1990prefix; martin2017parallelizing; smith2022simplified). It has been shown in (ladner1980parallel) that there exist algorithms for parallel scans when $\oplus$ is associative, i.e., $(a\oplus b)\oplus c=a\oplus(b\oplus c)$. Given a binary associative operator $\oplus$ and a sequence of $n$ elements $a_{1:n}$, the scan operation is defined as:
@@ 4. Experiments
In this section, we present our experimental setup, and the results to demonstrate the efficacy and efficiency of our proposed model against several state-of-the-art baselines. We conduct extensive experiments to answer the following research questions: RQ1: How does RecBLR perform against various sequential recommendation baselines? RQ2: What is the efficiency and scalability performance of RecBLR? RQ3: How different components of RecBLR contribute to its recommendation performance and efficiency? RQ4: How do different hyperparameter settings affect RecBLR?
@@ 4.1. Experimental Setup
Datasets. We choose the following public datasets for experiments: MovieLens-1M (harper2015movielens) is a widely-used dataset consists of 1 million movie rating. Gowalla (cho2011friendship) contains location check-ins from users of the Gowalla location-based social network. We also included Steam (kang2018self), which is is crawled from an online video game distribution platform. Amazon-Beauty and Amazon-Sports  (mcauley2015image) contain product ratings and reviews for the Beauty and Sports categories from Amazon. XLong (ren2019lifelong) is dataset from the Alibaba e-commerce platform with long user behavior sequences. We sample 5000 users from the original XLong dataset and use the reduced dataset for assessing efficiency and scalability. We create a user interaction sequence for each user by chronologically sorting their interaction behaviors based on timestamps. Following the approach adopted in previous studies (kang2018self), we exclude users and items with fewer than five recorded interactions. Table 2 summarizes the statistics of the datasets after preprocessing.
@@
Baselines. We compare the performance of RecBLR with several sequential recommendation baselines: FPMC (rendle2010factorizing) combines Markov chains and matrix factorization to provide better recommendations. Caser (tang2018personalized) employs horizontal and vertical convolutions to learn user preferences and sequential patterns. GRU4Rec (hidasi2015session) utilizes Gated Recurrent Units (chung2014empirical) to model user interactions and predict the next likely items of interest. NARM (li2017neural) is an RNN-based model with the attention mechanism using local and global encoders. SASRec (kang2018self) is a unidirectional Transformer-based model that leverages self-attention to adaptively capture user preferences from their behavior sequences. BERT4Rec (sun2019bert4rec) is a bidirectional Transformer-based model using Transformer encoders to model user behavior sequences. FMLP-Rec (zhou2022filter) is a recommendation model based on multilayer perceptron with learnable filters using Fast Fourier Transform. LRURec (yue2024linear) is an RNN-based recommendation model with complex-valued Linear Recurrent Units (orvieto2023resurrecting).
Evaluation Metrics. To assess the recommendation performance of all models, we employ three common metrics: Hit Ratio (HR), Normalized Discounted Cumulative Gain (NDCG), and Mean Reciprocal Rank (MRR). For each metric, we consider the top-$K$ recommendation with a cutoff at $K$, denoted as HR@$K$, NDCG@$K$, and MRR@$K$. We evaluate these metrics at $K=10$ and $K=20$.
Implementation Details. Our default model architecture utilizes two behavior modeling layers. We optimize the model using the Adam optimizer (kingma2014adam) with a learning rate of 0.001. A batch size of 2048 is used for training, and 4096 for evaluation. For the XLong dataset specifically, we use a smaller batch size of 512 for training and 1024 for evaluation. Training stops early if validation NDCG@10 fails to improve for 10 consecutive epochs. All models leverage an embedding dimension of 64. The maximum sequence length is set based on the average user interaction length: 1000 for XLong, 200 for MovieLens-1M, 100 for Gowalla, and 50 for other datasets. We refer to a commonly used recommendation library RecBole (zhao2021recbole) for further implementation details. Our baseline models are also implemented and evaluated based on RecBole to make a fair comparison.
4.2. Overall Performance (RQ1)
@@
  • RecBLR outperforms various baselines across all metrics and datasets. This demonstrates that our design significantly improves recommendation accuracy for users, regardless of the data platform.
@@
4.3. Efficiency and Scalability Study (RQ2)
@@
4.4. Ablation Study (RQ3)
To understand how different components contribute to RecBLR’s performance, we conduct an ablation study. We evaluate RecBLR variants with the following modifications: Single User Modeling Layer: Reduced the user modeling layers from two to one. BD-LRU Only: Replaced the gated recurrent layer with a basic Behavior-Dependent Linear Recurrent Unit (BD-LRU), lacking linear projections, gating mechanisms, and causal convolutions. Without Conv1D: Removed the unidirectional causal Conv1D layer from the gated recurrent layer. Without PFFN: Removed the position-wise feedforward network (PFFN). Without Custon Initialization: Replaced the custom LRU parameter initialization with random initialization. Table 3 shows the performance comparison between the default RecBLR and its variants. We have the following observations and conclusions: Generalizability: The default RecBLR achieves the best performance across most datasets, demonstrating strong generalization capabilities. Number of User Modeling Layers: While a single user modeling layer maintains decent performance, there is a trade-off between the number of layers and computational efficiency. A single layer can significantly improve training and inference speed. Importance of Gated Recurrent Layer: The BD-LRU-only variant, removing linear projections and gating mechanisms in the gated recurrent layer, demonstrates the importance of these components in enhancing the BD-LRU’s user behavior modeling capabilities. Impact of Conv1D: Removing the Conv1D layer, which captures temporal dependencies in sequential data, leads to slight performance degradation on most datasets except the Steam dataset, which may be due to the limited size and distribution of Steam. Impact of PFFN: The PFFN contributes to behavior modeling regardless of the underlying operator (LRU or attention mechanisms) in the behavior modeling layer. Removing the PFFN consistently reduces performance. LRU Initialization: The custom initialization scheme commonly used in LRUs  (orvieto2023resurrecting), plays a crucial role in enhancing the performance, especially for larger datasets like MovieLens-1M. Using random initialization yields lower performance.
@@ 4.5. Hyperparameter Sensitivity (RQ4)
Dropout Rate. We have experimented with various dropout rates across different datasets (Figure 6). We observe that selecting a higher dropout rate yields better performance for sparser datasets. For instance, the MovieLens-1M dataset, exhibiting relatively high sparsity, achieves optimal performance with a dropout rate of 0.2. Similarly, a dropout rate of 0.4 yields the best results for the Gowalla and Steam datasets. The sparse Amazon-Beauty and Amazon-Sports datasets achieve the highest performance at a dropout rate of 0.5. These observations suggest that considering the dataset’s sparsity is crucial when selecting dropout rates.
@@ 5. Related Work
In this section, we introduce the sequential recommendation task and discuss efficient sequence modeling approaches. We also highlight the promising role of efficient sequence modeling approaches in advancing the field of recommendation systems.
5.1. Sequential Recommendation
Sequential recommendation has undergone rapid advancements in recent years. Early approaches Markov Chains (MCs) with matrix factorization (rendle2010factorizing) to model user preferences based on historical interactions. Recently, deep learning methods have significantly enhanced the expressivity of sequential recommendation systems. Earlier neural approaches adopting Convolutional Neural Networks (tang2018personalized) and Recurrent Neural Networks (RNNs) (hidasi2015session; li2017neural) pioneer the application of neural network in sequence recommendation. Transformer-based architectures have become dominant due to their self-attention mechanism, which captures complex item relationships and improves accuracy. Additionally, the parallelizable nature of attention addresses the efficiency concerns of RNNs. While Transformers offer high accuracy, research continues to optimize inference efficiency. Alternative approaches, such as MLP-based models like FMLP-Rec (zhou2022filter) and LRURec (yue2024linear) based on linear recurrent units (LRUs), offer a balance between accuracy and efficiency. In our work, we further enhance the accuracy and efficiency of LRU-based recommendation models by utilizing real-valued behavior-dependent LRUs and hardware-efficient algorithms.
5.2. Efficient Sequence Modeling
To address the dilemma of performance, training efficiency, and inference efficiency in sequence modeling, various approaches have been proposed such as linear attention, linear RNNs, and state space models (SSMs). Recently, linear RNNs and SSMs have achieved both efficiency and effectiveness. SSMs (e.g., S4 (gu2021efficiently), S5 (smith2022simplified), Mamba (gu2023mamba)) have been widely adopted as alternatives to traditional RNNs, CNNs, and Transformers in general sequence modeling tasks. Inspired by the success of SSMs, Linear Recurrent Units  (orvieto2023resurrecting) reinvent RNNs, using linearization, diagonalization, stable exponential parameterization, and custom initialization to create RNN-based models with both strong performance and good efficiency. Within the recommendation domain, LRURec (yue2024linear) pioneers the application of linear recurrent units. However, there remains significant potential for exploring and adapting efficient sequence modeling approaches for sequential recommendation tasks. In our approach, we explore adapting LRUs for sequential recommendation. We employ real-valued LRUs with gating mechanisms to enhance their awareness and dependencies on user behavior. This enables our model to capture user behavior and preferences more effectively in recommendation tasks.

### v2

@@ Abstract.
Recommender Systems, Sequential Recommendation
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679717isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systems
@@ 1. Introduction
Sequential recommender systems are crucial for personalized online services, capturing evolving user preferences through past interactions and predicting future actions by effectively modeling sequential dependencies within historical user behaviors (chen2018sequential; xie2022contrastive; liu2024mamba4rec; lin2024rella). The core of sequential recommendation is the user behavior modeling techniques, which have evolved significantly over time, advancing from traditional methods such as Markov chains (he2016fusing; rendle2010factorizing) to sophisticated deep learning methods (hidasi2015session; yan2019cosrec). As illustrated in Figure 1, there is a triangle of golden principles for designing the user behavior modeling operators for sequential recommendation, i.e., training efficiency, low-cost inference, and strong performance.
@@
We choose the linear recurrent unit (LRU) (martin2017parallelizing; orvieto2023resurrecting) as the basic operator, and start by investigating its potential drawbacks in the case of sequential recommendation scenarios when considering the triangle of three golden principles. Firstly, the LRU operator is independent to the input user behaviors, i.e., every single user behavior is transformed through the same network, thereby contributing equally to the user preference representation. However, the diverse behaviors in the user sequence should make different contributions due to the multifaceted nature of user interests. Hence, it is essential to build a behavior-dependent operator to improve the performance of user behavior modeling. Secondly, previous work, i.e., LRURec (yue2024linear), simply follows the classical design of LRU by leveraging the eigenvalue decomposition in the complex domain. Nevertheless, as demonstrated in (gu2023mamba), complex elements are beneficial for continuous modalities like video and audio, but not for discrete sequence modeling like user behaviors. Moreover, complex-domain LRUs introduce a larger number of parameters compared to real-valued counterparts, potentially leading to overfitting and degenerated performance. Lastly, while the above two points both aim at the “strong performance” principle, they can largely hurt the training and inference efficiency due to the complicated design of behavior modeling operators (i.e., behavior-dependent and non-complex). Therefore, it is crucial to move beyond the simple model design, and consider hardware-aware parallelism acceleration algorithm for the rest two principles of “training efficiency” and “low-cost inference” simultaneously.
@@ 3.3. Hardware-Aware Parallel Scan
Since deep learning accelerators like GPUs are only optimized for operators like matrix multiplications and convolutions, we develop an efficient custom CUDA kernel for our BD-LRUs, implemented in OpenAI’s Triton (tillet2019triton), which is a language and compiler for GPU programming and parallel computing. Our hardware-aware algorithm is based on scans (blelloch1990prefix; martin2017parallelizing; smith2022simplified). It has been shown in (ladner1980parallel) that there exist algorithms for parallel scans when $\oplus$ is associative, i.e., $(a\oplus b)\oplus c=a\oplus(b\oplus c)$. Given a binary associative operator $\oplus$ and a sequence of $n$ elements $a_{1:n}$, the scan operation is defined as:
@@ 4.1. Experimental Setup
Datasets. We choose the following public datasets for experiments: MovieLens-1M (harper2015movielens) is a widely-used dataset consisting of 1 million movie ratings. Gowalla (cho2011friendship) contains location check-ins from users of the Gowalla location-based social network. Steam (kang2018self) contains reviews from an online video game distribution platform. Amazon-Beauty and Amazon-Sports  (mcauley2015image) contain product ratings and reviews for the Beauty and Sports categories from Amazon. XLong (ren2019lifelong) is a dataset from the Alibaba e-commerce platform with long user behavior sequences. We sample 5000 users from the original XLong dataset and use the reduced dataset for assessing efficiency and scalability. We create a user interaction sequence for each user by chronologically sorting their interaction behaviors based on timestamps. Following the approach adopted in previous studies (kang2018self), we exclude users and items with fewer than five recorded interactions. Table 2 summarizes the statistics of the datasets after preprocessing.
@@
Baselines. FPMC (rendle2010factorizing) combines Markov chains and matrix factorization for next-item prediction. Caser (tang2018personalized) employs horizontal and vertical convolutions to learn user preferences. GRU4Rec (hidasi2015session) utilizes Gated Recurrent Units (chung2014empirical) to model user interactions and predict the next-item. NARM (li2017neural) is an RNN-based model with the attention mechanism using local and global encoders. SASRec (kang2018self) is a unidirectional Transformer-based model leveraging self-attention to adaptively capture user preferences from their behavior sequences. BERT4Rec (sun2019bert4rec) uses bidirectional Transformer encoders to model user behavior sequences. FMLP-Rec (zhou2022filter) is based on multilayer perceptron with learnable filters using Fast Fourier Transform. LRURec (yue2024linear) is an RNN-based recommendation model with complex-valued Linear Recurrent Units (orvieto2023resurrecting).
Evaluation Metrics. We employ three common metrics: Hit Ratio (HR), Normalized Discounted Cumulative Gain (NDCG), and Mean Reciprocal Rank (MRR). For each metric, we consider the top-$K$ recommendation with a cutoff at $K$, denoted as HR@$K$, NDCG@$K$, and MRR@$K$. We evaluate these metrics at $K=10$ and $K=20$.
Implementation Details. Our default model architecture utilizes two behavior modeling layers. We adopt Adam optimizer (kingma2014adam) with a learning rate of 0.001. A batch size of 2048 is used for training, and 4096 for evaluation. For the XLong dataset specifically, we use a smaller batch size of 512 for training and 1024 for evaluation. Training stops early if validation NDCG@10 fails to improve for 10 consecutive epochs. All models leverage an embedding dimension of 64. The maximum sequence length is set based on the average user interaction length: 1000 for XLong, 200 for MovieLens-1M, 100 for Gowalla, and 50 for other datasets. We refer to a commonly used recommendation library RecBole (zhao2021recbole) for further implementation details and baseline comparisons. Our code is available at https://github.com/chengkai-liu/RecBLR.
4.2. Overall Performance
@@
  • RecBLR outperforms all baselines for all metrics and datasets. This demonstrates that our design significantly improves recommendation accuracy for users, regardless of the data platform.
@@
4.3. Efficiency and Scalability Study
@@
4.4. Ablation Study
To understand how different components contribute to RecBLR’s performance, we conduct an ablation study. We evaluate RecBLR variants with the following modifications: Single Behavior Modeling Layer: Reduced the behavior modeling layers from two to one. BD-LRU Only: Replaced the gated recurrent layer with a basic Behavior-Dependent Linear Recurrent Unit (BD-LRU), lacking linear projections, gating mechanisms, and causal convolutions. Without Conv1D: Removed the unidirectional causal Conv1D layer from the gated recurrent layer. Without PFFN: Removed the position-wise feedforward network (PFFN). Without Custon Initialization: Replaced the custom LRU parameter initialization with random initialization. Table 3 shows the performance comparison between the default RecBLR and its variants. We have the following observations and conclusions: Generalizability: The default RecBLR achieves the best performance across most datasets, demonstrating strong generalization capabilities. Number of Behavior Modeling Layers: While a single behavior modeling layer maintains decent performance, there is a trade-off between the number of layers and computational efficiency. A single layer can significantly improve training and inference speed. Importance of Gated Recurrent Layer: The BD-LRU-only variant, removing linear projections and gating mechanisms in the gated recurrent layer, demonstrates the importance of these components in enhancing the BD-LRU’s user behavior modeling capabilities. Impact of Conv1D: Removing the Conv1D layer, which captures temporal dependencies in sequential data, leads to slight performance degradation on most datasets except the Steam dataset, which may be due to the limited size and distribution of Steam. Impact of PFFN: The PFFN contributes to behavior modeling regardless of the underlying operator (LRU or attention mechanisms) in the behavior modeling layer. Removing the PFFN consistently reduces performance. LRU Initialization: The custom initialization scheme commonly used in LRUs  (orvieto2023resurrecting), plays a crucial role in enhancing the performance, especially for larger datasets like MovieLens-1M. Using random initialization yields lower performance.
@@ 4.5. Hyperparameter Sensitivity
Dropout Rate. We have experimented with various dropout rates across different datasets (Figure 6). We observe that selecting a higher dropout rate yields better performance for sparser datasets. For instance, the MovieLens-1M dataset, exhibiting relatively low sparsity, achieves optimal performance with a dropout rate of 0.2. Similarly, a dropout rate of 0.4 yields the best results for the Gowalla and Steam datasets. The sparse Amazon-Beauty and Amazon-Sports datasets achieve the highest performance at a dropout rate of 0.5. These observations suggest that considering the dataset’s sparsity is crucial when selecting dropout rates.
@@ 5. Related Work
Sequential Recommendation. Sequential recommendation has undergone rapid advancements in recent years. Early approaches Markov Chains (MCs) with matrix factorization (rendle2010factorizing) to model user preferences based on historical interactions. Recently, deep learning methods have significantly enhanced the expressivity of sequential recommendation systems (xi2023bird; lin2024can; lin2023map; xi2023towards). Earlier neural approaches adopting Convolutional Neural Networks (tang2018personalized) and Recurrent Neural Networks (RNNs) (hidasi2015session; li2017neural) pioneer the application of neural network in sequence recommendation. Transformer-based architectures have become dominant due to their self-attention mechanism, which captures complex item relationships and improves accuracy. Additionally, the parallelizable nature of attention addresses the efficiency concerns of RNNs. While Transformers offer high accuracy, research continues to optimize inference efficiency. Alternative approaches, such as MLP-based models like FMLP-Rec (zhou2022filter) and LRURec (yue2024linear) based on linear recurrent units (LRUs), offer a balance between accuracy and efficiency.
Efficient Sequence Modeling. To address the dilemma of performance, training efficiency, and inference efficiency in sequence modeling, various approaches have been proposed such as linear attention, linear RNNs, and state space models (SSMs). Recently, linear RNNs and SSMs have achieved both efficiency and effectiveness. SSMs (e.g., S4 (gu2021efficiently), S5 (smith2022simplified), Mamba (gu2023mamba)) have been widely adopted as alternatives to traditional RNNs, CNNs, and Transformers in general sequence modeling tasks. Inspired by the success of SSMs, Linear Recurrent Units  (orvieto2023resurrecting) reinvent RNNs, using linearization, diagonalization, stable exponential parameterization, and custom initialization to create RNN-based models with both strong performance and good efficiency. Within the recommendation domain, LRURec (yue2024linear) pioneers the application of linear recurrent units. However, there remains significant potential for exploring and adapting efficient sequence modeling approaches for sequential recommendation tasks.